# Meta-Generalized-Gradient Approximation made Magnetic

### v1

@@ Abstract
In Table 1 we list the spin magnetic moments (calculated analytically in the Gaussian basis set) and the magnetic energies (in meV), calculated as energy differences w.r.t. the self-consistent spin-restricted solution for the ferromagnetic metals Fe, Co and Ni. All calculations employ the experimental geometries reported in Ref. Tran et al., 2020, unless explicitly stated otherwise. Comparisons are provided against the LDA (employing the parametrization of the Ceperly-Alder correlation energies of the uniform gas by Vosko, Wilk and Nusair, and the usual Slater-Dirac energy functional for exchange SVWN5),Vosko et al. (1980) the GGA of Perdew, Burke and Ernzerhof (PBE),Perdew et al. (1996) and the original spin-unrestricted SCAN of Sun, Ruzsinszky and Perdew.Sun et al. (2015) Comparison against experimental data for the spin-magnetic moments are also available in the last row of the table.
@@
In accord with previous reports,Fu and Singh (2018); Tran et al. (2020); Mejía-Rodríguez and Trickey (2019) we find an agreement with experiments at the level of LDA, which is not significantly changed with the GGA. SCAN, on the other hand, displays the previously reported overmagnetization in all ferromagnetic metals, both in its significant overestimation of the magnetic moments, as well as overestimation of the magnetic energies.Fu and Singh (2018); Tran et al. (2020); Mejía-Rodríguez and Trickey (2019) In the case of Fe, for example, the self-consistently predicted magnetic moment of 2.62 $\mu_{b}$ by SCAN represents an overestimation of around 30 %. Overmagnetization of around 10 % and 15 % w.r.t. experimental values are predicted by SCAN for Co and Ni respectively, while the LDA and GGA both fall within the experimental uncertainty. Magnetic energies of SCAN are consistently more than double the values predicted by the LDA and GGA for the ferromagnetic metals. The problem is completely solved by switching to our gauge-invariant mSCAN approach, which provides similar magnetic moments (and magnetic energies) as the LDA and GGA on Co and Ni, while the predicted magnetic moment on Fe of 2.15 $\mu_{b}$ is the closest to the experimental range of 1.98-2.08 $\mu_{b}$.
Table 2: Same as in Table 1 but for non-magnetic metals. The calculations employed a ferromagnetic guess density matrix and the spin-moments were relaxed along the self-consistent field procedure.
V Pd
LDA 0.00 (0.00) 0.00 (0.00)
PBE 0.00 (0.00) 0.20 (0.70)
SCAN 0.33 (1.23) 0.44 (17.2)
mSCAN 0.00 (0.00) 0.00 (0.00)
EXP 0.00Weber et al. (1997) 0.00Kudasov and Korshunov (2007)
@@
Table 3: Spin magnetization in $\mu_{b}$ and, within parenthesis, band gaps (in eV) of antiferromagnetic Mott insulators.
@@
00footnotetext: We remove the orbital contributions provided in Ref. Tran et al., 2020 from the experimental total (spin+orbital) magnetic moments (see caption of Table VI of Ref. Tran et al., 2020).
The improvement provided by our gauge-invariant mSCAN approach is even more striking when applied to the non-magnetic materials V and Pd (see Table 2). We perform the calculations starting from a ferromagnetic guess and let the spin moments relax during the self-consistent procedure. Like in previous works,Fu and Singh (2018); Tran et al. (2020) SCAN calculations incorrectly predict a net magnetic moment of 0.33 and 0.44 $\mu_{b}$ on V and Pd respectively. In the case of Pd, the incorrect magnetic behaviour is also predicted by the GGA of PBE, even though the correct non-magnetic behaviour is predicted by the simpler LDA. Remarkably, the correct non-magnetic behaviour of V and Pd is predicted by the proposed gauge-invariant mSCAN approach.
Although the failures of SCAN on itinerant magnets are daunting, accurate magnetic moments and gaps have been predicted in the antiferromagnetic Mott insulators FeO, CoO and NiO, which exhibit more localized magnetism.Zhang et al. (2020); Tran et al. (2020) Values obtained from our mSCAN on these highly-correlated Mott insulators are provided in Table 3, including a comparison against the experiment both on magnetic moments and fundamental band gaps (in parenthesis, given in eV). The LDA and GGA incorrectly predict a metallic behaviour for FeO and CoO, while both SCAN and mSCAN provide the qualitatively correct insulating behaviour. All functionals provide similar values for the spin magnetic moments in the Mott insulators, although both moments and band gaps are larger with SCAN.
@@
We conclude by investigating the performance of mSCAN on structural optimizations in all tested systems (see Table 4). A previous attempt at partially fixing the spin moments of SCAN by varying the $c_{1x}$ parameter has produced significantly worsened results for lattice constants, with errors increased by around 50%.Tran et al. (2020) Here, mean absolute errors (MAE) decrease in the order LDA, GGA, mSCAN (SCAN), with values of 0.914, 0.682, 0.380 (0.365), which is again consistent with the principle of the Jacob’s ladder of DFT.
Table 4: Lattice constants of all systems as obtained from geometry optimizations with different DFAs. Square brackets denote $\mathbf{c}$ lattice parameters and round brackets denote lattice angles, when not fixed by symmetry. Mean absolute errors (MAE) w.r.t. the experiments (see Ref. Tran et al., 2020) are also reported.
Fe Co Ni V Pd FeO CoO NiO MAE
LDA 2.760 2.429 [3.917] 3.353 2.905 3.769 5.624 (28.323) 5.487 (28.559) 4.930 (33.958) 0.914
PBE 2.845 2.490 [4.024] 3.425 2.977 3.840 5.685 (29.310) 5.515 (29.824) 5.102 (33.747) 0.682
SCAN 2.860 2.462 [3.994] 3.389 2.956 3.801 5.607 (29.605) 5.151 (33.707) 5.076 (33.662) 0.365
mSCAN 2.785 2.454 [3.956] 3.387 2.956 3.799 5.576 (29.655) 5.160 (33.308) 5.053 (33.698) 0.380
EXP 2.867 2.507 [4.070] 3.523 3.028 3.881 5.576 (33.557) 5.210 (33.557) 5.108 (33.557) -
Conclusions.  In this work, the soundness of the third rung of the Jacob’s ladder of density functional theory (DFT) has been established for magnetic systems by a magnetic modification of the strongly constrained and appropriately normed semilocal density functional – mSCAN. mSCAN puts aside the constraint of zero correlation energy for any one-electron spin-polarized density (i.e. in mSCAN, $\widetilde{\alpha}$ does not vanish for single-particle orbitals). Moreover, the exchange component is not obtained from the spin-scaling relations [compare Eq. (14a) with Eq. (8)]. Crucially, mSCAN derives from a more general framework — namely, spin-current-DFT (instead of collinear spin DFT) — which dictates invariance of the xc energy functional under (local) conjoint U(1)$\times$×SU(2) gauge transformations as a most important exact constraint. Even though the mSCAN is subsequently specialized to the collinear case, the use of a more general framework has exposed an important dependence of the functional on the gradient of the collinear spin density – a dependence that would have been missed in the conventional approach.
While the question of including all the relevant exact constraints will have to be explored further — leading to possible further improvements — the present work shows that mSCAN resolves the most daunting issues that have plagued the application of the original SCAN to prototypical magnetic materials.

### v2

@@ Abstract
Next, we consider also materials. In Table 1 we list the spin magnetic moments (calculated analytically in the Gaussian basis set) and the magnetic energies (in meV), calculated as energy differences w.r.t. the self-consistent spin-restricted solution for the FM metals Fe, Co and Ni. All calculations employ the experimental geometries reported in Ref. Tran et al., 2020, unless explicitly stated otherwise. Comparisons are provided against the LDA (employing the parametrization of correlation energies by Vosko, Wilk and Nusair, and the usual Slater-Dirac energy functional for exchange SVWN5),Vosko et al. (1980) the GGA of PBE,Perdew et al. (1996) and the original spin-unrestricted SCAN.Sun et al. (2015) Comparison against experimental data for the spin-magnetic moments are also given.
@@
In accord with previous reports,Fu and Singh (2018); Tran et al. (2020); Mejía-Rodríguez and Trickey (2019) we find an agreement with experiments at the level of LDA, which is not significantly changed with the GGA. SCAN displays the previously reported overmagnetization in all FM metals, both in its significant overestimation of the magnetic moments, as well as overestimation of the magnetic energies.Fu and Singh (2018); Tran et al. (2020); Mejía-Rodríguez and Trickey (2019) In the case of Fe, for example, the self-consistently predicted magnetic moment of 2.62 $\mu_{b}$ by SCAN represents an overestimation of around 30 %. Overmagnetization of around 10 % and 15 % w.r.t. experimental values are predicted by SCAN for Co and Ni respectively, while the LDA and GGA both fall within the experimental uncertainty. Magnetic energies of SCAN are consistently more than double the values predicted by the LDA and GGA for the FM metals. The problem is completely solved by switching to our gauge-invariant mSCAN approach, which provides similar magnetic moments (and magnetic energies) as the LDA and GGA on Co and Ni, while the predicted magnetic moment on Fe of 2.15 $\mu_{b}$ is the closest to the experimental range of 1.98-2.08 $\mu_{b}$.
The improvement provided by our approach is even more striking when applied to the non-magnetic metals V and Pd (see Table 2), as well as the difficult case of Cr2.Larsson et al. (2022); Zhang et al. (2020a) We perform the calculations starting from a spin-polarized guess and let the spin moments relax during the self-consistent procedure. Like in previous works,Fu and Singh (2018); Tran et al. (2020) SCAN calculations incorrectly predict a net magnetic moment of 0.33 and 0.44 $\mu_{b}$ on V and Pd respectively. In the case of Pd, the incorrect magnetic behaviour is also predicted by PBE. For Cr2, the LDA, PBE and SCAN all predict a spurious magnetization at the equilibrium geometry (ground state properties are reported in Table S1 of the SM). Remarkably, the correct non-magnetic behaviour of V, Pd and Cr2 is predicted by the proposed gauge-invariant mSCAN approach. As for itinerant two-dimensional systems, in Graphene, SCAN incorrectly predicts a magnetic insulating phase,Zhang et al. (2020a) while mSCAN correctly provides a non-magnetic, metallic state, with a Dirac point at the zone corners (Table S2).
Table 2: Same as in Table 1 but for non-magnetic systems.
V Pd Cr2
LDA 0.00 (0.00) 0.00 (0.00) 2.69 (0.30)
PBE 0.00 (0.00) 0.20 (0.70) 5.32 (0.87)
SCAN 0.57 (6.27) 0.44 (17.2) 5.62 (10.3)
mSCAN 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
EXP 0.00Weber et al. (1997) 0.00Kudasov and Korshunov (2007) 0.00Larsson et al. (2022)
00footnotetext: Values in eV.
Although the failures of SCAN on itinerant magnets are daunting, accurate magnetic moments and gaps have been predicted in the AFM Mott insulators FeO, CoO and NiO, which exhibit more localized magnetism.Zhang et al. (2020b); Tran et al. (2020) Values obtained from our mSCAN on these highly-correlated Mott insulators are provided in Table 3, including a comparison against the experiment both on magnetic moments and fundamental band gaps (in parenthesis, given in eV). The LDA and GGA incorrectly predict a metallic behaviour for FeO and CoO, while both SCAN and mSCAN provide the qualitatively correct insulating behaviour. All functionals provide similar values for the spin magnetic moments in the Mott insulators, although both moments and band gaps are larger with SCAN. We also consider a topological AFM – the Chern magnet TbMn6Sn6, in which calculation of spin moments is a reported challenge for SDFT.Annaberdiyev et al. (2023) With mSCAN, we obtain excellent agreement against the experimental values (with large improvement over SCAN and PBE) in Table S3.
@@ V Pd Cr2
Table 3: Spin moments (in $\mu_{b}$/atom) and, within parenthesis, band gaps (in eV) of AFM Mott insulators.
@@
00footnotetext: We remove the orbital contributions provided in Ref. Tran et al., 2020 from the experimental total moments (see caption of Table VI of Ref. Tran et al., 2020).
@@
Next, we investigate the performance of mSCAN on structural optimizations in all tested systems (see Table S4). A previous attempt at partially fixing the spin moments of SCAN by varying the $c_{1x}$ parameter has produced significantly worsened results for lattice constants, with errors increased by around 50%.Tran et al. (2020) Here, mean absolute errors (MAE) decrease in the order LDA, GGA, mSCAN (SCAN), with values of 0.663, 0.521, 0.295 (0.319), which is again consistent with the principle of the Jacob’s ladder of DFT.
Let us now scrutinize the dissociation of first-row diatomic molecules involving open-shell atoms.not (g) Fig. S3 reports the case of H2 compared to the accurate result of Ref. Kolos and Wolniewicz, 1968: the mSCAN and SCAN energies match up until an intermolecular separation of about 3 Å, at which point a spurious magnetization appears in the spin-unrestricted LDA, PBE and SCAN solutions. mSCAN provides a correct vanishing spin magnetization up to about 3.5 Å. In the C2 molecule LDA, PBE and SCAN provide spurious magnetization even at the equilibrium geometry, which is resolved by mSCAN, see Fig. S4 and Table S5. mSCAN also provides an improvement over SCAN, with respect to the reference total energies of Ref. Booth et al., 2011 on C2, both at the equilibrium geometry and at large distances. Fig. S5 reports the dissociation of the N2 molecule. mSCAN provides a significant improvement over LDA, PBE and SCAN at distances larger than about 1.5 Å, where strong correlation becomes important (as also corroborated by comparison of multi- and single-reference coupled-cluster treatments).Li and Paldus (2008)
Figure 1: Magnetization in $\mu_{b}$ of the Cr3 molecule with the nc SCAN functional.
Noncollinear extension.  Fully self-consistent SCDFT MGGA calculations for non-collinear magnetic states are reported in a parallel work.Huebsch et al. (2025) Ref. Huebsch et al., 2025 also calls for a gauge-invariant nc-extension of SCAN.
Thus, we finally solve Eq. (10) using the nc forms in Eq.s (13a) and (13b). Technical details are provided in the SM, where we also demonstrate that the functional provides a non-zero xc magnetic torque, which vanishes over the full space. Last but not least, spin-orbit coupling can be accounted for self-consistently. So, our proposal fulfills the most pressing desiderata for a nc approach.
We applied it to the Cr3 (bond length of 3.7 Bohr), which has been an important model system in previous studies.Scalmani and Frisch (2012); Desmarais et al. (2024a); Huebsch et al. (2025); Tancogne-Dejean et al. (2023); Pu et al. (2023) The magnetization reported in Fig. 1 is indicative of a spin-frustrated system. We observe a spin-moment of 1.38 $\mu_{B}$, which is significantly lower than the SCDFT-MGGA value of 3.15 $\mu_{B}$ reported in Ref. Huebsch et al., 2025, but closer to the LDA value of 1.68 $\mu_{B}$.
Conclusions.  While it remains a challenge to assess whether a semi-local density functional may be derived to embrace all the relevant non-magnetic and magnetic (both collinear and non-collinearly polarized) exact constraints, the core findings of this work solves most urgent failures for magnetism and resumes progress towards improved density functional approximations.